OpenStudy (anonymous):

I am so confused.....PLEASE HELP I understand part c but not a and b. A right triangle has base (x - 7) units and height (2x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. Part B: What is the area of the triangle? Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials.

OpenStudy (anonymous):

Please help, I'm lost>

OpenStudy (anonymous):

Use a part of Pythagorean theorem for A. The equation will be: (x - 7)^2 + (2x - 10)^2 Do not square root the answer since it is telling you to get a square of the hypotenuse. For part B, use a formula for an area of a triangle. (b*h) / 2 = (x - 7)(2x - 10) / 2

OpenStudy (anonymous):

Oh thank you. That is what I di for Part a but I squared it. So, I will unsquare it now.

OpenStudy (anonymous):

So for part b, the answer is what you wrote? Because otherwise how would I do the problem (x-7)(2x-10)/2 ? @cindykim95

OpenStudy (anonymous):

(x-7)(2x-10) = 2x^2 - 24x + 70 (2x^2 - 24x + 70)/2 = x^2 - 12x + 35

OpenStudy (anonymous):

Thank you so much. I always have so much trouble with math. Can you help with another one? @cindykim95

OpenStudy (anonymous):

You're welcome :) And yes, I can help you.

OpenStudy (anonymous):

The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n2 + 2,500n - 4,000 Part A: What are the zeroes of the above function and what do they represent? Show your work. Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work. Part C: What is the axis of symmetry of the function P(n)?

OpenStudy (anonymous):

Please wait for a second. I'm working on my paper now.

OpenStudy (anonymous):

Do you know what is the derivative by any chance? I mean, have you ever learned about that concept?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

Part A: I don't understand the question. Sorry.. :( Seriously, I've never seen this kind of question in my life. Can you explain this to me if you can? Part B: p(n)' (Derivative of p(n) = -500n + 2500 Let p(n)' so it becomes 0 = -500n + 2500 -2500 = -500n n = 5. If you put 5 into the original function, it becomes -250(5^2) + 2500(5) - 4000 = 2250 (This is the final answer) Part C: There is a formula for getting the axis of symmetry for a quadratic equation. For ax^2+bx+c = 0, -b/2a. In this equation, b = 2500 and a = -250 If you plug the number, -2500/(2*-250) = 5 (this is the axis of symmetry)

OpenStudy (anonymous):

@cindykim95, that was all that was written in the question

OpenStudy (anonymous):

In case you want to know what the derivative is, for example, Derivative of x^5 = 5x^4 Derivative of 5x^4 = 20x^3 Derivative of 20x^3 = 60x^2 Derivative of 60x^2 = 120x Derivative of 120x is 120 Derivative of 120 is 0. Do you kind of see the pattern here?

OpenStudy (anonymous):

Yes, I think so

OpenStudy (anonymous):

I think in part a the zeros represent the ticket @cindykim95

OpenStudy (anonymous):

@Cindykim95... thank you so much. You really helped me out a lot! When I see it, I understand it more. I just can't figure it out. :) Thanks again.

OpenStudy (anonymous):

You're welcome :)